OpenStudy (jamanorthi):

The light sensitive substance in black and white photography film is AgBr. photons provide the energy necessary to transfer an electron from Br- to Ag+ to produce Ag and Br. a) if a minimum energy of 2.00×10^5J/mol is needed for this process, what is the minimum energy needed by each photon?

OpenStudy (sunnnystrong):

Energy States: \[En=E _{0}(\frac{ Z^2 }{ n^2 })\] En=energy state in joules n=shell # Z=atomic # Eo=ground state of Hydrogen electron (-2.178*10^(-18) J/n^(2)) \[\lambda = \frac{ hc }{ \Delta E }\] Delta E= Ef-Ei h=Planck's Constant (6.6262*10^(-34) m^(2)/s c= speed of light (2.998*10^(8) m/s) \[c=\lambda \nu \] c=speed of light lambda=wave length nu=frequency